OpenStudy (anonymous):

What are the domain and range of the function Y = -2 |x + 1|+ 4???? How do you find the range and domain??? I need help!

OpenStudy (mr.math):

The domain of an absolute value function is all real numbers (i.e \((-\infty,\infty)\).

OpenStudy (mr.math):

The range of an absolute value function is all non-negative value, that's \[|x+1|\ge0 \implies -2|x+1|\le0 \implies -2|x+1|+4\le 4\] So, the range is \((-\infty,4]\).

OpenStudy (anonymous):

Ok thank you :)

OpenStudy (mr.math):

You're welcome.